OpenStudy (anonymous):

sum1 help Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8)

jimthompson5910 (jim_thompson5910):

First find the slope m = (y2-y1)/(x2-x1) m = ( 8- 5)/(6-4) I'll let you finish to find the slope. Tell me what you get.

OpenStudy (anonymous):

m=3-2 m=1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( 8- 5)/(6-4) m = 3/(6-4) m = 3/2

jimthompson5910 (jim_thompson5910):

so the slope is 3/2

OpenStudy (anonymous):

oh so the bttm doesnt change

jimthompson5910 (jim_thompson5910):

not sure what you mean, the bottom goes from 6-4 to 2, so it does change

OpenStudy (anonymous):

oh so it has to = a fraction

jimthompson5910 (jim_thompson5910):

yes, you'll often get a fraction for the slope

jimthompson5910 (jim_thompson5910):

not always though

OpenStudy (anonymous):

so u reduce it right

jimthompson5910 (jim_thompson5910):

3/2 is fully reduced

OpenStudy (anonymous):

ik but im saying like if it wer 2/6 u can reduce it

jimthompson5910 (jim_thompson5910):

yes, 2/6 would reduce to 1/3

OpenStudy (anonymous):

oh wait look at my answers all 4 say 3/2?

jimthompson5910 (jim_thompson5910):

now that you know the slope is 3/2, we know m = 3/2 one of the points is (4,5), so x = 4 and y = 5 y = mx+b 5 = (3/2)*(4) + b 5 = (3/2)*(4/1) + b 5 = (3*4)/(2*1) + b 5 = 12/2 + b 5 = 6+b Keep going to find the value of b

OpenStudy (anonymous):

imma attach it

jimthompson5910 (jim_thompson5910):

what do you get when you solve 5 = 6+b for b?

OpenStudy (anonymous):

wait wat do i write for my answer bc they all say 3/2 x +or- something

jimthompson5910 (jim_thompson5910):

what do you get when you solve 5 = 6+b for b? ignore the answer choices for now

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

so we know that m = 3/2 and b = -1 which means \[\Large y = mx+b\] becomes \[\Large y = \frac{3}{2}x-1\]

OpenStudy (anonymous):

ok